OpenStudy (anonymous):

Solve using quadratic formula: 10x^2+10x=-3 I got my answer as -10+/- i(sq.root) 5 --------- 10 is that right?

OpenStudy (anonymous):

\[10x^2+10x+3=0\] \[a=10,b=10,c=3\] ok i think you forgot to divide -10 also. it should be \[\frac{-5+\sqrt{5}i}{10}\]

OpenStudy (anonymous):

i will write it out if you like

OpenStudy (anonymous):

Okay yeah that might help.

OpenStudy (anonymous):

\[x=\frac{-10+\sqrt{10^2-4\times 10 \times 3}}{2\times 10}\] \[=\frac{-10+\sqrt{100-120}}{20}\] \[=\frac{-10+\sqrt{-20}}{20}\]

OpenStudy (anonymous):

so far so good. \[=\frac{-10+\sqrt{4\times 5\times -1}}{20}\] \[=\frac{-10+2\sqrt{5}i}{20}\] \[=\frac{2(-5+\sqrt{5}i}{20})\] \[=\frac{-5+\sqrt{5}i}{10}\] think this is right.

OpenStudy (anonymous):

of course you should have the 'plus or minus"